### v1

# Rho-1: Not All Tokens Are What You Need

Abstract
Previous language model pre-training methods have uniformly applied a next-token prediction loss to all training tokens. Challenging this norm, we posit that “Not all tokens in a corpus are equally important for language model training”. Our initial analysis delves into token-level training dynamics of language model, revealing distinct loss patterns for different tokens. Leveraging these insights, we introduce a new language model called Rho-1. Unlike traditional LMs that learn to predict every next token in a corpus, Rho-1 employs Selective Language Modeling (SLM), which selectively trains on useful tokens that aligned with the desired distribution. This approach involves scoring pretraining tokens using a reference model, and then training the language model with a focused loss on tokens with higher excess loss. When continual pretraining on 15B OpenWebMath corpus, Rho-1 yields an absolute improvement in few-shot accuracy of up to 30% in 9 math tasks. After fine-tuning, Rho-1-1B and 7B achieved state-of-the-art results of 40.6% and 51.8% on MATH dataset, respectively — matching DeepSeekMath with only 3% of the pretraining tokens. Furthermore, when pretraining on 80B general tokens, Rho-1 achieves 6.8% average enhancement across 15 diverse tasks, increasing both efficiency and performance of the language model pre-training.
Figure 1: We continual pretrain 1B and 7B LMs with 15B OpenWebMath tokens. Rho-1 is trained with our proposed Selective Language Modeling (SLM), while baselines are trained using causal language modeling. SLM improves average few-shot accuracy on GSM8k and MATH by over 16%, achieving the baseline performance 5-10x faster.
Figure 2: Upper: Even an extensively filtered pretraining corpus contains token-level noise. Left: Previous Causal Language Modeling (CLM) trains on all tokens. Right: Our proposed Selective Language Modeling (SLM) selectively applies loss on those useful and clean tokens.
1 Introduction
Scaling up model parameters and dataset size has consistently elevated the next-token prediction accuracy in large language models, yielding significant advancements in artificial intelligence (Kaplan et al., 2020; Brown et al., 2020; OpenAI, 2023; Team et al., 2023). However, training on all available data is not always optimal or feasible. As a result, the practice of data filtering has become crucial, using various heuristics and classifiers (Brown et al., 2020; Wenzek et al., 2019) to select training documents. These techniques significantly improve data quality and boost model performance.
However, despite thorough document-level filtering, high-quality datasets still contain many noisy tokens that can negatively affect training, as illustrated in Figure 2 (Upper). Removing such tokens might alter the text’s meaning, while overly strict filtering could exclude useful data (Welbl et al., 2021; Muennighoff et al., 2024) and lead to biases (Dodge et al., 2021; Longpre et al., 2023). Furthermore, research indicates that the distribution of web data does not inherently align with the ideal distribution for downstream applications (Tay et al., 2022; Wettig et al., 2023). For example, common corpus at the token level may include undesirable content like hallucinations or highly ambiguous tokens that are hard to predict. Applying the same loss to all tokens can result in wasted computation on non-beneficial tokens, possibly limiting LLM’s potential to merely mediocre intelligence.
To explore how language models learn at the token level, we initially examined training dynamics, particularly how the token-level loss evolves during usual pretraining. In §2.1, we evaluated the model’s token perplexity at different checkpoints and categorized tokens into different types. Our findings reveal that significant loss reduction is limited to a select group of tokens during training. Many tokens are “easy tokens” that are already learned, and some are “hard tokens” that exhibit variable losses and resist convergence. These tokens can lead to numerous ineffective gradient updates.
Based on these analyses, we introduce Rho-1 models trained with a novel Selective Language Modeling (SLM) objective. As shown in Figure 2 (Right), this approach inputs the full sequence into the model and selectively removes the loss of undesired tokens. The detailed pipeline is depicted in Figure 4: First, SLM trains a reference language model on high-quality corpora. This model establishes utility metrics to score tokens according to the desired distribution, naturally filtering out unclean and irrelevant tokens. Second, SLM uses the reference model to score each token in a corpus using its loss (§2.2). Finally, we train a language model only on those tokens that exhibit a high excess loss between the reference and the training model, selectively learning the tokens that best benefit downstream applications (§2.2).
We show through comprehensive experiments that SLM significantly enhances token efficiency during pretraining and improves performance on downstream tasks. Furthermore, our findings indicate that SLM effectively identifies tokens relevant to the target distribution, resulting in improved perplexity scores on benchmarks for models trained with the selected tokens. §3.2 shows the effectiveness of SLM on math continual pretraining: both 1B and 7B Rho-1 outperform CLM-trained baselines by over 16% on the GSM8k and MATH datasets. SLM reaches baseline accuracy up to 10x faster, as shown in Figure 1. Remarkably, Rho-1-7B matches the state-of-the-art performance of DeepSeekMath-7B using only 15B tokens, compared to the 500B tokens required by DeepSeekMath. Upon fine-tuning, Rho-1-1B and 7B achieve 40.6% and 51.8% on MATH, respectively. Notably, Rho-1-1B is the first 1B LM to exceed 40% accuracy, nearing the early GPT-4’s CoT performance of 42.5%. §3.3 confirms the efficacy of SLM in general pretraining: Training Tinyllama-1B on 80B tokens with SLM improves 6.8% on average across 15 benchmarks, with gains over 10% in code and math tasks.
Figure 3: The loss of four categories of tokens during pretraining. (a) shows the loss of H→H, L→H, H→L, and L→L tokens during pretraining. (b) and (c) show three cases of fluctuating tokens’ loss in L→L and H→H during pretraining, respectively.
2 Selective Language Modeling
2.1 Not All Tokens Are Equal: Training Dynamics of Token Loss
Our investigation begins with a critical look at how individual tokens’ losses evolve during standard pre-training. We continue pre-training Tinyllama-1B with 15B tokens from OpenWebMath, saving checkpoints after every 1B tokens. We then evaluate token-level loss at these intervals using the validation set of approximately 320,000 tokens. Figure 3(a) reveals a striking pattern: tokens fall into four categories based on their loss trajectory—persistent high loss (H→H), increasing loss (L→H), decreasing loss (H→L), and consistent low loss (L→L). For further details on these categories, see §B.1. Our analysis uncovers that a mere 26% of tokens show a notable loss reduction (H→L), while the majority (51%) remain in the L→L category, indicating they have already been learned. Interestingly, 11% of the tokens are persistently challenging (H→H), likely due to high aleatoric uncertainty (Hüllermeier and Waegeman, 2021). Additionally, 12% of tokens experience an unexpected loss increase (L→H) during training.
Our second observation is that a significant number of token losses exhibit persistent fluctuations, and resist convergence. The loss of many L→L and H→H tokens, as depicted in Figure 3 (b) and (c), show high variance during training. In §B.2, we visualize and analyze the content of these tokens and find that many of them are noisy, which is consistent with our hypothesis.
Consequently, we learn that the loss associated with each token during training does not decrease smoothly like the overall loss; instead, there is a complex training dynamic among different tokens. If we can select the appropriate tokens for the model to focus on during training, we may be able to stabilize the trajectory of the model’s training and enhance its efficiency.
2.2 Selective Language Modeling
Overview
Inspired by the practice of reference model in document-level filtering, we propose a simple pipeline of token-level data selection, termed “Selective Language Modeling (SLM)”. Our method comprises three steps, as depicted in Figure 4. We begin by training a reference model on a curated, high-quality dataset. This model then assesses the loss of each token within the pretraining corpus. In the final phase, we train the language model selectively, focusing on tokens with high excess loss between the training and reference model. The intuition is that tokens with high excess loss are more learnable and better aligned with the desired distribution, naturally excluding tokens that are either irrelevant or of low quality. Below, we provide a detailed description of each step.
Reference Modeling
We begin by curating a high-quality dataset that reflects the desired data distribution. We train a reference model (RM) using standard cross-entropy loss on the curated data. The resulting RM is then used to assess the token loss within a larger pretraining corpus. We compute the reference loss ($\mathcal{L}_{\text{ref}}$) of a token $x_{i}$ based on the probability that the RM assigns to this token. The calculation is formalized as follows:
$\mathcal{L}_{\text{ref}}(x_{i})=-\log P(x_{i}|x{<i})$ (1)
By evaluating $\mathcal{L}_{\text{ref}}$ for each token, we establish the reference loss for selective pretraining, allowing us to focus on the most influential tokens in language modeling.
Figure 4: The pipeline of Selective Language Modeling (SLM). SLM optimizes language model performance by concentrating on valuable, clean tokens during pre-training. It involves three steps: (Step 1) Initially, train a reference model on high-quality data. (Step 2) Then, score each token’s loss in a corpus using the reference model. (Step 3) Finally, train the language model selectively on tokens that show higher excess loss compared to the reference loss.
Selective Pretraining
Note that causal language modeling (CLM) employs the cross-entropy loss:
$\mathcal{L}_{\text{CLM}}(\theta)=-\frac{1}{N}\sum_{i=1}^{N}\log P(x_{i}|x_{<i};\theta)$ (2)
Here, $\mathcal{L_{\text{CLM}}}(\theta)$ represents the loss function parameterized by model $\theta$. $N$ is the length of the sequence, $x_{i}$ is the $i$-th token in the sequence, and $x_{<i}$ represents all tokens before the $i$-th token. In contrast, Selective Language Modeling (SLM) trains the language model with a focus on tokens that exhibit a high excess loss when compared to the reference model. The excess loss ($\mathcal{L}_{\Delta}$) for a token $x_{i}$ is defined as the difference between the current training model loss ($\mathcal{L}_{\theta}$) and the reference loss:
$\mathcal{L}_{\Delta}(x_{i})=\mathcal{L}_{\theta}(x_{i})-\mathcal{L}_{\text{ref}}(x_{i})$ (3)
We introduce a token selection ratio $k\%$, which determines the proportion of tokens to be included based on their excess loss. The cross-entropy loss for the selected tokens is computed as follows:
$\mathcal{L}_{\text{SLM}}(\theta)=-\frac{1}{N*k\%}\sum_{i=1}^{N}I_{k\%}(x_{i})\cdot\log P(x_{i}|x_{<i};\theta)$ (4)
Here, $N*k\%$ defines the number of tokens that fall within the top $k\%$ of excess loss. The indicator function $I_{k\%}(x_{i})$ is defined as:
$I_{k\%}(x_{i})=\begin{cases}1&\text{if }x_{i}\text{ is in the top }k\%\text{ of }\mathcal{L}_{\Delta}\\ 0&\text{otherwise}\end{cases}$ (5)
This ensures that the loss is applied only to the tokens that are deemed most beneficial for the language model to learn from. In practice, token selection can be implemented by ranking the tokens in a batch according to their excess loss and using only the top $k\%$ of tokens for training. This process eliminates the loss for undesired tokens without incurring additional costs during pretraining, making our approach both efficient and easily integrated.
3 Experiments
We continually pretrained models in both mathematical and general domain and designed ablation and analysis experiments to understand the effectiveness of SLM.
3.1 Experimental Setup
Reference Model Training
To train our mathematical reference model, we gathered a dataset of 0.5B high-quality, math-related tokens. This dataset is a blend of synthetic data from GPT (Yu et al., 2024; Huang et al., 2024) and manually curated data (Yue et al., 2024; Ni et al., 2024). For the general reference model, we compiled a corpus of 1.9B tokens from open-source datasets, such as Tulu-v2 (Ivison et al., 2023) and OpenHermes-2.5 (Teknium, 2023). We trained the reference models for 3 epochs. The maximum learning rate was set at 5e-5 for 1B models and 1e-5 for 7B models, applying a cosine decay schedule. We set the maximum sequence lengths to 2048 for 1B models and 4096 for 7B models, packing multiple samples into these lengths for model input. In all main experiments, we initialized the continual pretraining model and the reference model with the same base model.
Table 1: Few-shot CoT reasoning results of math pretraining. All models are tested with few-shot prompting. Previous best results are highlighted in blue, while our best results are in purple. $*{}^{*}$nly unique math-related tokens are calculated. For Rho-1, we calculate only the selected tokens that are used for training. ${}^{\dagger}$e use OpenAI’s MATH subset (Lightman et al., 2023) for evaluation, since some original test samples have been used in public training sets such as PRM800k. ${}^{\ddagger}$he SAT only has 32 four-choice problems, so we average our results over the last three checkpoints, if available.
Model $|\bm{\theta}|$ Data
Uniq.
Train
Toks
GSM8K ${}^{\dagger}$ SVAMP ASDiv MAWPS TAB MQA
MMLU
STEM
${}^{\ddagger}$ AVG
1-2B Base Models
Tinyllama 1.1B - - - 2.9 3.2 11.0 18.1 20.4 12.5 14.6 16.1 21.9 13.4
Phi-1.5 1.3B - - - 32.4 4.2 43.4 53.1 66.2 24.4 14.3 21.8 18.8 31.0
Qwen1.5 1.8B - - - 36.1 6.8 48.5 63.6 79.0 29.2 25.1 31.3 40.6 40.0
Gemma 2.0B - - - 18.8 11.4 38.0 56.6 72.5 36.9 26.8 34.4 50.0 38.4
DeepSeekLLM 1.3B OWM 14B 150B 11.5 8.9 - - - - - 29.6 31.3 -
DeepSeekMath 1.3B - 120B 150B 23.8 13.6 - - - - - 33.1 56.3 -
Continual Pretraining on Tinyllama-1B
Tinyllama-CT 1.1B OWM 14B 15B 6.4 2.4 21.7 36.7 47.7 17.9 13.9 23.0 25.0 21.6
Rho-1-Math 1.1B OWM 14B 9B 29.8 14.0 49.2 61.4 79.8 25.8 30.4 24.7 28.1 38.1
$\Delta$40% +23.4 +11.6 +27.5 +24.7 +32.1 +7.9 +16.5 +1.7 +3.1 +16.5
Rho-1-Math 1.1B OWM 14B 30B 36.2 15.6 52.1 67.0 83.9 29.0 32.5 23.3 28.1 40.9
$\geq$ 7B Base Models
LLaMA-2 7B - - 14.0 3.6 39.5 51.7 63.5 30.9 12.4 32.7 34.4 31.4
Mistral 7B - - 41.2 11.6 64.7 68.5 87.5 52.9 33.0 49.5 59.4 52.0
Minerva 8B - 39B 164B 16.2 14.1 - - - - - 35.6 - -
Minerva 62B - 39B 109B 52.4 27.6 - - - - - 53.9 - -
Minerva 540B - 39B 26B 58.8 33.6 - - - - - 63.9 - -
LLemma 7B PPile 55B 200B 38.8 17.2 56.1 69.1 82.4 48.7 41.0 45.4 59.4 50.9
LLemma 34B PPile 55B 50B 54.2 23.0 67.9 75.7 90.1 57.0 49.8 54.7 68.8 60.1
Intern-Math 7B - 31B 125B 41.8 14.4 61.6 66.8 83.7 50.0 57.3 24.8 37.5 48.7
Intern-Math 20B - 31B 125B 65.4 30.0 75.7 79.3 94.0 50.9 38.5 53.1 71.9 62.1
DeepSeekMath 7B - 120B 500B 64.1 34.2 74.0 83.9 92.4 63.4 62.4 56.4 84.4 68.4
Continual Pretraining on Mistral-7B
Mistral-CT 7B OWM 14B 15B 42.9 22.2 68.6 71.0 86.1 45.1 47.7 52.6 65.6 55.8
Rho-1-Math 7B OWM 14B 10.5B 66.9 31.0 77.8 79.0 93.9 49.9 58.7 54.6 84.4 66.2
$\Delta$30% +24.0 +8.8 +9.2 +8.0 +7.8 +4.8 +11.0 +2.0 +18.8 +10.4
Pretraining Corpus
For mathematical reasoning, we utilize the OpenWebMath (OWM) dataset (Paster et al., 2023), which comprises approximately 14B tokens sourced from math-related web pages in the Common Crawl. In the general domain, we combine the SlimPajama (Daria et al., 2023) and StarCoderData (Li et al., 2023a) (both part of the Tinyllama corpus) with OpenWebMath, training on a total of 80 billion tokens with a mix ratio of 6:3:1.
Pretraining Setting
For math pretraining, we continue pretraining on the Tinyllama-1.1B model (Zhang et al., 2024) and the Mistral-7B model (Jiang et al., 2023) with learning rates of 8e-5 and 2e-5, respectively. For general domain, we set the learning rate for Tinyllama-1.1B model to 1e-4. The batch size is uniformly set to 1M tokens for both domains. Regarding the token selection ratio, we use 60% for the Tinyllama-1.1B model and 70% for the Mistral-7B model.
Baseline Setting
We use models that have been continually pretrained (Tinyllama-CT and Mistral-CT) through regular causal language modeling as baselines. Moreover, we compare Rho-1 with well-known and top-performing baselines, including Gemma (Team et al., 2024), Qwen1.5 (Bai et al., 2023), Phi-1.5 (Li et al., 2023b), DeepSeekLLM (DeepSeek-AI, 2024), DeepSeekMath (Shao et al., 2024a), CodeLlama (Roziere et al., 2023), Mistral (Jiang et al., 2023), Minerva (Lewkowycz et al., 2022), Tinyllama (Zhang et al., 2024), LLemma (Azerbayev et al., 2023), and InternLM2-Math (Ying et al., 2024). For fine-tuning results, we also compare with previous best models MAmmoTH(Yue et al., 2024) and ToRA(Gou et al., 2024).
Evaluation Setup
To comprehensively evaluate pretrained models, we compare their few-shot capabilities and fine-tuning performance across a variety of tasks. We adopt the lm-eval-harness [^1] https://github.com/EleutherAI/lm-evaluation-harness for general tasks, and math-eval-harness [^2] https://github.com/ZubinGou/math-evaluation-harness for math tasks. We use vllm (v0.3.2) (Kwon et al., 2023) to speed up inference. Further details on our evaluation can be found in Appendix C.
Table 2: Tool-integrated reasoning results of math pretraining.
Model Size Tools SFT Data GSM8k MATH SVAMP ASDiv MAWPS TAB GSM-H AVG
Used for SFT?
Previous Models
GPT4-0314 - - 92.0 42.5 93.1 91.3 97.6 67.1 64.7 78.3
GPT4-0314 (PAL) - - 94.2 51.8 94.8 92.6 97.7 95.9 77.6 86.4
MAmmoTH 70B MI-260k 76.9 41.8 82.4 - - - - -
ToRA 7B ToRA-69k 68.8 40.1 68.2 73.9 88.8 42.4 54.6 62.4
ToRA 70B ToRA-69k 84.3 49.7 82.7 86.8 93.8 74.0 67.2 76.9
DeepSeekMath 7B ToRA-69k 79.8 52.0 80.1 87.1 93.8 85.8 63.1 77.4
Our Pretrained Models
TinyLlama-CT 1B ToRA-69k 51.4 38.4 53.4 66.7 81.7 20.5 42.8 50.7
Rho-1-Math 1B ToRA-69k 59.4 40.6 60.7 74.2 88.6 26.7 48.1 56.9
$\Delta$ +8.0 +2.2 +7.3 +7.5 +6.9 +6.2 +5.3 +6.2
Mistral-CT 7B ToRA-69k 77.5 48.4 76.9 83.8 93.4 67.5 60.4 72.6
Rho-1-Math 7B ToRA-69k 81.3 51.8 80.8 85.5 94.5 70.1 63.1 75.3
$\Delta$ +3.8 +3.4 +3.9 +1.7 +1.1 +2.6 +2.7 +2.7
3.2 Math Pre-training Results
Few-shot CoT Reasoning Results
We evalute base models prompting with few-shot chain-of-thought (CoT) (Wei et al., 2022a) examples following previous works (Lewkowycz et al., 2022; Azerbayev et al., 2023; Shao et al., 2024b). As results shown in Table 1, in comparison to continue pretraining directly, Rho-1-Math has achieved the average few-shot accuracy improvement of 16.5% on 1B models and 10.4% on 7B models. Furthermore, after training for multiple epochs on OpenWebMath, we find that Rho-1 could further increase the average few-shot accuracy to 40.9%. Compared to DeepSeekMath-7B, which pretrained on 500 billion math-related tokens, Rho-1-7B pretrained on only 15 billion tokens (selecting 10.5 billion tokens) achieved comparable results, demonstrating the efficiency of our approach.
Tool-Integrated Reasoning Results
We fine-tune Rho-1 and baseline models on 69k ToRA corpus (Gou et al., 2024), consisting of 16k GPT-4-generated trajectories in a tool-integrated reasoning format, and 53k answer-augmented samples using LLaMA. As presented in Table 2, Rho-1-1B and Rho-1-7B achieved a state-of-the-art 40.6% and 51.8% on MATH dataset, respectively. On some unseen tasks (e.g., TabMWP and GSM-Hard), Rho-1 also demonstrates a certain degree of generalizability, with an average few-shot accuracy improvement of 6.2% on the Rho-1-Math-1B and 2.7% on Rho-1-Math-7B.
Figure 5: General pretraining results. We continual pretraining Tinyllama-1B on 80G general tokens. Tinyllama-CT is etrained with CLM, while Rho-1 is trained with our proposed SLM.
3.3 General Pre-training Results
We confirm the efficacy of the SLM in general pretraining by continual training Tinyllama-1.1B on 80 billion tokens. The results depicted in Figure 5 indicate that although Tinyllama has already undergone extensive training on the majority of these tokens, the application of SLM yields an average enhancement of 6.8% across 15 benchmarks compared to direct continual pretraining. The improvements were especially pronounced in code and math tasks, exceeding 10%.
Figure 6: The dynamics of pretraining loss and downstream loss. (a) and (c) represent the loss of tokens selected/unselected by SLM during pretraining in both SLM and CLM methods, while (b) represents the loss of the SLM and CLM methods on downstream corpora. We tested the above results through the process of pretraining with a total of 4 billion tokens.
3.4 Analysis
Selected Token Loss Aligns Better with Downstream Performance
We use the reference model to filter tokens and explore the changes in validation loss after training on all/selected tokens, while observing their relationship with downstream loss. As shown in Figure 6, we pretrain approximately 4B tokens and displayed the variation curves of loss on different pretraining method and validation sets during the pretraining process. We can observe that on the tokens selected by the reference model, the decrease in average loss of the Rho-1 is more significant compared to regular pretraining. On the contrary, on the unselected tokens, the decrease in average loss of the regular pretraining is more significant. If we associate fig (a), fig(b) with fig(c), it is not difficult to find that the model trained on the selected tokens has a more significant decrease in downstream loss, while ordinary pretraining, although reducing the average loss of all tokens during the training phase, is difficult to have a significant decrease in downstream loss. Therefore, we expect that selecting tokens for pretraining is more efficient.
Figure 7: The relationship between the selected tokens / unselected tokens loss in SLM and downstream task performance. The y-axis represents the average few-shot accuracy on GSM8k and MATH. The x-axis represents the average loss on selected tokens / unselected tokens at corresponding checkpoint(2B, 5B, 8B, 11B, and 14B).
Moreover, We relate the selected tokens’ loss to its downstream task performance via a power law in Figure 7, which is similar to a concurrent study (Gadre et al., 2024). Observing the curve fitted from the data points in the graph, the average loss of tokens selected by the SLM shows a positive correlation with the performance of downstream tasks, whereas the average loss of tokens not selected exhibits a negative correlation with downstream task performance. Therefore, it is not necessary for the all tokens’ loss to decrease to benefit the model’s ultimate performance. See Appendix D for more details.
[Uncaptioned image] Figure 8: The PPL of tokens selected by different checkpoint. We test the PPL of the tokens selected at 2B, 5B, 8B, 11B, and 14B.
[Uncaptioned image] Figure 9: Effect of token select ratio. We train 1B LM with SLM objective on 5B tokens.
What Tokens are Selected with SLM?
We aim to analyze the tokens selected by the SLM method in pretraining to further explore its working mechanism. To this end, we visualize the token selection process during the training of Rho-1 using the OpenWebMath. In §E.1, we have highlighted in blue the tokens that were retained during actual pretraining. We observe that the majority of tokens chosen by the SLM method are closely related to mathematics, effectively training the model on the parts of the original corpus that are pertinent to mathematical content.
Furthermore, we investigated the differences in token filtering across various checkpoints during the training process and tested the perplexity of these tokens on different checkpoints. As illustrated in Figure 8, we found that the tokens selected by later checkpoints tend to have higher perplexity towards the later stages of training and lower perplexity in the earlier stages. This may suggest that the model first optimizes tokens with a larger learnable space, thereby increasing learning efficiency. Moreover, we noticed a sample-wise “double descent” (Nakkiran et al., 2021) on the loss of selected tokens, where the select token’s perplexity initially increases before decreases. This might be an effect of selecting tokens based on excess loss, targeting those most in need at each checkpoint.
Effect of Token Select Ratio
We investigate the impact of token selecting ratios of the SLM. Generally, the selecting ratio is defined by heuristic rules, similar to the approach previously employed in the training of Masked Language Models (MLMs) (Devlin et al., 2019; Liu et al., 2019). As shown in Figure 9, the selected tokens is suitable for accounting for about 60% of the original tokens.
Weak-to-Strong Generization
Table 3: Weak-to-Strong generization result on math benchmark.
Model GSM8K ${}^{\dagger}$ SVAMP ASDiv MAWPS TAB MQA
MMLU
STEM
${}^{\ddagger}$ AVG
Llama-2-7B-CT 28.4 13.6 50.3 62.8 79.5 37.6 34.1 41.6 43.5 43.5
Llama-2-7B-CT w/ 1B RM 29.8 16.0 55.5 63.7 80.4 37.9 34.3 38.2 43.8 44.4
Apart from the main experiments where we use the same base model for the reference and continual pretraining, we also investigate if a smaller reference model can effectively guide the pretraining of a larger model. We use Tinyllma-1.1B as reference model and continual pretraining Llama-2-7B on math. Results presented in Table 3 indicate that, despite the considerable gap between the small and large models (Li et al., 2023c), employing the small reference model to token selection can still yield benefits to the pre-training of the larger model. If reference and training models have different vocabularies, one can consider performing token alignment (Wan et al., 2024; Fu et al., 2023), which we leave for future work.
4 Related Works
Pretraining Data Optimization
The objective of optimizing pre-training corpora is to maximize the performance and efficiency of language model training by improving the quality and scale of the pretrain data mixture. This typically includes data collecting through crawling (Raffel et al., 2020) or synthesis (Polu and Sutskever, 2020; Gunasekar et al., 2023), de-duplication (Lee et al., 2021; Kandpal et al., 2022; Tirumala et al., 2023), filtering and selection (Albalak et al., 2024), as well as data composition (Xie et al., 2024) and curriculum (Chen et al., 2024; MA et al., 2024).
Data Selection
Data selection for fine-tuning has been extensively studied, focusing on improving quality (Li et al., 2023d), diversity (Liu et al., 2024), and distribution matching (Li et al., 2023e; Xia et al., 2024; Ni et al., 2024). For pretraining, various lightweight filters are utilized (Albalak et al., 2024), including heuristic-based (e.g., language and item count filtering), classifier-based (Brown et al., 2020), and perplexity-based approaches (Wenzek et al., 2019). The massive public RedPajama-Data-v2 dataset (Computer, 2023), for example, leverages over 40 quality indicators for data filtering and reweighting. Nevertheless, strict filtering like blocklist (Raffel et al., 2020) and Safety API filtering (Welbl et al., 2021), have been found to hurt evaluation loss or induce bias (Dodge et al., 2021). To our knowledge, we are the first to explore token-level data selection, aimed at enhancing data quality and information density at the most fundamental granularity.
Language Model Training Dynamics
Investigating the training dynamics of language models is essential for understanding their behavior throughout the training process. This research includes studying internal representations (Saphra and Lopez, 2018), the acquisition of linguistic knowledge (Choshen et al., 2021; Liu et al., 2021), and the phenomenon of grokking (Power et al., 2022). The analysis by Xia et al. (2022) is the most related to ours, which examines token-level training trajectories in models of varying sizes. Our findings, however, diverge from those of Xia et al. (2022), who posit that tokens with little change in perplexity are “already learned”. We identify a spectrum of token patterns, including “easy tokens” and “hard tokens” that resist convergence. Recognizing this, we propose a method of selective language modeling that targets the influential tokens, optimizing the learning process.
Scaling Laws
Scaling laws guide us in discovering the impact of factors such as parameter count, data size, and compute on language model performance and behavior. These studies usually focus on predicable scaling though power law (Kaplan et al., 2020; Hernandez et al., 2021), optimal resource allocation (Hoffmann et al., 2022), downstream tasks (Wei et al., 2022b; Isik et al., 2024; Gadre et al., 2024), architectures (Tay et al., 2022), memorization (Tirumala et al., 2022; Carlini et al., 2022; Henighan et al., 2023; Biderman et al., 2024), and repeating data (Hernandez et al., 2022; Muennighoff et al., 2024; Xue et al., 2024). Most scaling laws on model performance study cross-entory loss on all training tokens, while we focus on the tokens loss of desired distributions.
5 Discussion and Future Work
Generalization
In math continual pretraining, as depicted in Figure 6, training exclusively with SLM leads to quickly convergence to the domain focused by the reference model, accompanied by a significant rise in the loss of unselected tokens. Although no adverse effects, like biases, have been observed from the increased loss yet, a general pretraining loss on text and code may prevent overfitting (Goodhart and Goodhart, 1984), as suggested by Ouyang et al. (2022) and Azerbayev et al. (2023). Furthermore, future efforts could broaden the corpus scope of the reference model, and enlarge the pretraining data size, as exemplified by DeepSpeedMath (Shao et al., 2024a).
Scalability
Due to budget constraints, we have only verified the effectiveness of our method on smaller models (<=7B parameters) and smaller datasets (<100B tokens). Smaller models benefit significantly from removing the loss of irrelevant tokens and focusing on important ones. However, it’s possible that very large models trained on extensive corpora may naturally develop this inductive bias to compress useful data (i.e., compressing everything), although it may sounds inefficient for now. Therefore, future works should study whether this selective language modeling technique can scale to very large models and data (Kaplan et al., 2020).
Is training a reference model necessary?
To score tokens, we need a high-quality reference model. This could be a base model trained with a small amount of high-quality data, or a performant open-source model. In fact, since we only need input logprobs or perplexity from reference model, we could even utilize more powerful proprietary model APIs. We can input tokens and use the log probabilities of the input returned by the API as reference scores. We leave this for future works.
How to improve upon SLM?
There are many natural extensions of SLM, e.g., reweighting tokens instead of selecting may improve robustness; using a reference model as a reward model to guide pretraining with reinforcement learning; adopting multiple reference models to reduce overfitting; designing token-level curriculum learning and iterative strategies for continuous improvements, etc.
Expanding the Use of SLM
SLM may be extended to supervised fine-tuning to address the noise and distribution mismatches in many SFT datasets. Another potential application is alignment, e.g., by training a reference model to emphasize helpfulness, truthfulness, and harmlessness, we may obtain a base model that is natively aligned during the pretraining stage.
Appendix A Author Contributions
Zhenghao Lin designed and implemented detailed token selection process, conducted extensive preliminary experiments, developed the pre-training and evaluation pipeline, conducted most of the pre-training experiments and analysis, implemented baselines, and significantly contributed to the writing. Zhibin Gou presented a preliminary proposal, introduced the method of using excess loss for reweighting tokens, compiled high-quality corpora, trained reference models, set up the fine-tuning and evaluation pipelines, designed the experimental analysis, and significantly contributed to the writing. Yeyun Gong proposed the initial project and co-led the project with Weizhu Chen, they offered extensive advice and guidance on experiments and writing, and oversaw team collaboration and resource management. Xiao Liu, Yelong Shen, Ruochen Xu, Chen Lin, Yujiu Yang, Jian Jiao, and Nan Duan offered research mentorship, coordinated the project, and contributed to the writing.
Appendix B Analysis and Visualization of Tokens in Pretraining
B.1 More Details of Four Categories Tokens
Figure 10: Sample text containing four categories of tokens. Among them, blue represents tokens of categorie H→L, green indicates tokens of categorie L→L, yellow signifies tokens of categorie H→H, and red denotes tokens of categorie L→H.
We categorize tokens into four categories: H→H, L→H, H→L, L→L. During the training process, we collected the loss of each token after training on each 1 billion tokens training data. We then used linear fitting and took the difference in loss between the first and last points as evidence of whether the loss decreased during the training process.
Specifically, suppose we have a sequence of token’s loss $(l_{0},l_{1},...,l_{n})$. Our goal is to minimize the sum of the squares of the differences between each data point and its linear predictive value:
$f(a,b)=\text{minimize}\sum_{i=0}^{n}(l_{i}-(ax_{i}+b))^{2},$ (6)
where $x_{0}=0$ is the initial checkpoint and $x_{n}=n$ is the final checkpoint. Substituting these into the fitted equation, we can obtain the Loss values at the start and end after fitting: $\mathcal{L}_{\text{start}}=b$ and $\mathcal{L}_{\text{end}}=an+b$. The change in loss can then be expressed as: $\Delta\mathcal{L}=\mathcal{L}_{\text{end}}-\mathcal{L}_{\text{start}}$. Meanwhile, we represent the average Loss of the last checkpoint as $\mathcal{L}_{\text{mean}}$.
Next, we can classify the tokens based on $\Delta\mathcal{L}$ and the $\mathcal{L}_{\text{mean}}$. We categorize tokens with $\Delta\mathcal{L}<-0.2$ as H→L (loss decreases from high to low) category tokens, and tokens with $\Delta\mathcal{L}>0.2$ as L→H (loss increases from low to high) category tokens. If $-0.2\leq\Delta\mathcal{L}\leq 0.2$ and $l_{n}\leq\mathcal{L}_{\text{mean}}$, then tokens are classified as L→L (loss remains low); if $l_{n}>\mathcal{L}_{\text{mean}}$, they are classified as H→H (loss remains high). In Figure 10, we visualize examples of the four categories of tokens in actual text.
B.2 Non-Converging Tokens in Pretrainig
Figure 11: An example of an abnormal state of token perplexity during pretrainig process. The tokens highlighted in orange represent tokens that were significant abnormalities during the pretraining process.
In §2.1, we mentioned that during the training process, only a minority of tokens belong to the H→L category. Among the remaining categories of H→H and L→L tokens, there are tokens that exhibit significant fluctuations during training. Furthermore, there are instances where H→L tokens are not effectively learned. Therefore, in our analysis, we specifically select those tokens from these categories that demonstrate considerable variability and distinct loss. We visualize these tokens that exhibit abnormal behavior during the training process. As illustrated in Figure 11, we find that the majority of these tokens originate from rather chaotic corpora. For instance, the corpora may include a mix of custom symbols, unintelligible gibberish, and information such as timetables and bibliographic references. Within a segment of normal text, there may also be fluctuations in the usage of common conjunctions, word suffixes, and punctuation marks. The latter may not necessarily be disastrous for training; in fact, it could represent a normal occurrence. However, if we can effectively mitigate the losses caused by the former, it might lead to more stable and efficient model training.
Appendix C Evalution Details
C.1 Math Evalution
We conducted a comprehensive evaluation of the model across various math reasoning benchmarks, encompassing a range of difficulties from elementary to university level, multiple mathematical domains, and diverse question types including multiple-choice and open-ended questions. Our benchmarks include GSM8k (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), GSM-Hard (Gao et al., 2022), SVAMP (Patel et al., 2021), ASDIV (Miao et al., 2020), MAWPS (Koncel-Kedziorski et al., 2016), TabMWP (TAB) (Lu et al., 2023), MathQA (MQA) (Amini et al., 2019), MMLU-STEM (Hendrycks et al., 2020), and SAT (Azerbayev et al., 2023).
C.2 General Evalution
In the evaluation of general domain, we followed the lm-evaluation-harness (Gao et al., 2023) and evalute model on MMLU (Hendrycks et al., 2020), BBH (Suzgun et al., 2022), AGIEval (Zhong et al., 2023), ARC-Easy and ARC-Challenge (Clark et al., 2018), BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), Hellaswag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), OpenBookQA (Mihaylov et al., 2018). On HumanEval (Zheng et al., 2023) and TydiQA (Clark et al., 2020), we follow the evaluation pipeline of open-instrcut (Ivison et al., 2023) and report Pass@1 and Pass@10 for HumanEval and F1 for TydiQA. For MBPP (Austin et al., 2021) benchmark, we follow the evaluation pipeline of DeepSeek-Coder (Guo et al., 2024), and report Pass@1 and Pass@10.
Appendix D Relate the Selected Tokens’ Loss to Downstream Task Performance
In this section, we declare the details about correlating the loss of selected tokens with the performance of downstream tasks. Concurrent study has explored similar methods to study the impact of scaling laws with the performance of models in downstream tasks (Gadre et al., 2024). Our analysis here differs in that it aims to elucidate the relationship between the decrease/increase in loss for selected/unselected tokens and the model’s performance on downstream tasks.
We use the average accuracy of MATH and GSM8K as the standard for measuring downstream tasks performance of model. Based on the trend of data points in Figure 7, we propose the relationship between the average accuracy of downstream tasks and selected/unselected tokens’ loss,
$Acc(\mathcal{L})=\log(a*\mathcal{L}+c)$ (7)
The parameters $a$ and $c$ are fitted from the data. If the loss of selected tokens $\mathcal{L}_{s}$ is used for fitting, then $a>0$. Conversely, if the loss of unselected tokens $\mathcal{L}_{us}$ is used for fitting, then $a<0$. Therefore, we believe that training the model on selected tokens can effectively improve its performance on downstream tasks, while unselected tokens may have a detrimental effect on the model’s performance in downstream tasks.
Appendix E Examples of Tokens Selected by SLM
E.1 Token Selected Examples
Figure 12: Specific examples of selecting tokens during the ToT training process of the Rho-1. The tokens marked in blue represent the actual tokens trained during the ToT training process, while the remaining black tokens are not trained during the ToT training process.
In Figure 12, we present several examples of tokens selected by the SLM method, with content marked in blue indicating the tokens actually chosen during the pretraining process.
E.2 Dynamic Token Selected
Figure 13: An example of dynamic token selection changes during the training process, which illustrated with five different score levels represented by deep blue, light blue, black, light orange, and dark orange. The bluer the color indicates a higher tendency for the token to be selected, while the more orange the color suggests a lower tendency for the token to be selected.
In Figure 13, we display the dynamic changes in token selection tendencies throughout the SLM training process. We chose four checkpoints during the training process (0%, 33%, 66%, and 100%) to analyze the current tendencies in token selection. The preferences for token selection are indicated by different colors, ranging from high to low preference, typically represented as deep blue, blue, black, orange, and dark orange, respectively.